A concave lens has a focal length of -36 cm. Find
Answer #1

6.

f = focal length = - 36 cm

do = object distance = 19 cm

di = image distance

using the lens equation

1/di + 1/do = 1/f

1/di + 1/19 = 1/(-36)

di = - 12.4 cm

m = magnification = -di/do = -(-12.4)/19 = 0.65